MsBrains (ms-brains):

PLEASE HELP!! I JUST NEED SOMEBODY TO CHECK MY ANSWERS! I WILL FAN AND MEDAL!!

OpenStudy (dn_20):

What answer ?

MsBrains (ms-brains):

I posted it...in the attached link

MsBrains (ms-brains):

@DN_20

OpenStudy (anonymous):

Multiply each side of the equation by 4/3 leaving a single x in the LHS. Simplify the RHS.

OpenStudy (dn_20):

So you substitute x = 12 in the equation \[\frac{ 3 }{ 4 }x = 9\]\[12\frac {3} {4} =9\]\[9 = 9\] to have x with the multiplication of 1, you multiply the number of its opposite, which is 4/3 \[\frac{4}{3}*\frac{3}{4} = 9*\frac{4}{3}\]\[x = 12\] So the answer is 4/3

MsBrains (ms-brains):

Yup. Thanks

OpenStudy (dn_20):

you're welcome :)